# Spatio-temporal Early Prediction based on Multi-objective Reinforcement Learning

### v1

@@ Abstract.
Accuracy and timeliness are indeed often conflicting goals in prediction tasks. Premature predictions may yield a higher rate of false alarms, whereas delaying predictions to gather more information can render them too late to be useful. In applications such as wildfires, crimes, and traffic jams, timely predictions are vital for safeguarding human life and property. Consequently, finding a balance between accuracy and timeliness is crucial. In this paper, we propose a spatio-temporal early prediction model based on Multi-Objective reinforcement learning that can either implement an optimal policy given a preference or infer the preference based on a small number of samples. The model addresses two primary challenges: 1) enhancing the accuracy of early predictions and 2) providing the optimal policy for determining the most suitable prediction time for each area. Our method demonstrates superior performance on three large-scale real-world datasets, surpassing existing methods in early spatio-temporal prediction tasks.
Spatio-temporal data, Graph neural network, Early prediction, Responsible AI
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computing methodologies Spatial and physical reasoningccs: Applied computing Forecasting
@@ 1. Introduction
Figure 1. Example of three methods for early spatio-temporal prediction. Each circle is a node (e.g. sensor) with recorded value (such as speed) over time. The blue plane represents the prediction time, and the recorded values after the blue plane are not used for prediction. The adaptive early prediction method adjusts data usage and dynamically determines the prediction time for different nodes. The node colour at time $T$ indicates the accuracy of prediction, green indicates that the predicted value matches the ground truth, and red indicates otherwise.
@@
Spatio-temporal early prediction focuses on anticipating future events by examining patterns that change over space and time as early as possible, which is essential to many real-world applications such as epidemiology, environmental studies, and public safety, where early detection can lead to better management or prevention outcomes. For instance, within the context of forecasting the spread of diseases, the timeliness of predictions is considerably more critical than their accuracy. Forecasts that are accurate but delayed can lead to the unnecessary loss of thousands of lives. On the other hand, a prediction that is approximately accurate but made in a timely manner can significantly reduce the impact, saving lives.
@@
Figure 1 shows that existing approaches to predicting spatio-temporal events are categorised into three distinct strategies: Traditional, Fixed, and Adaptive. For example, Hochreiter et al. (Hochreiter and Schmidhuber, 1997) propose a traditional prediction method that considers the entire recorded data before the prediction. However, this approach can be time-consuming, particularly with large datasets, and its requirement for complete prior data collection often results in delayed predictions, which is unsuitable for time-sensitive applications. Conversely, Li’s fixed early prediction approach operates within a set observation period (Li et al., 2017), facilitating timely forecasts within a specific window. Although this strategy enhances promptness, its fixed timing can result in inaccuracies if the observation window is not optimally selected. Moreover, its rigidity in adjusting to temporal data variations and the risk of selection bias in determining the observation window could compromise prediction precision. The third approach, named adaptive early prediction, dynamically adjusts the prediction time based on data characteristics. This idea is inspired by Hartvigsen, who applied this idea in the time-series prediction area (Hartvigsen et al., 2019). This innovative method aims to strike an optimal balance between prediction accuracy and timeliness, thereby attempting to optimise both the timeliness and accuracy of predictions. Its dynamic adaptability provides a significant edge in handling changes over time.
@@
Although Hartivigsen has tried to apply the concept of early prediction in time-series area (Hartvigsen et al., 2019) and achieved some good results in the medical diagnosis, many challenges remain, especially in the spatio-temporal prediction. (1) Real-time Bi-objective Balance in Highly Complex Environment: Spatio-temporal data demand dynamic solutions, being more complex and subject to rapid changes than time-series data. Traditional multi-objective optimisation methods, which are static and computationally intensive, fail to adapt efficiently across different environments. At the same time, early prediction necessitates immediate results, as even a one-minute delay can lead to unfavourable outcomes. (2) Spatio-temporal Comprehensive Dependency: Relying solely on distance correlation for spatio-temporal predictions may fall short in quickly capturing essential data characteristics. Although distance correlation effectively reflects spatial connections among nodes, it overlooks the crucial temporal aspects embedded within the data, which are vital for precise forecasting. (3) Hidden Preferences Discovery: The equilibrium between timeliness and accuracy varies across tasks, revealing diverse preferences. Identifying these subtle preferences is a complex challenge that demands a profound comprehension of the goals at hand and formulating an ideal balance between promptness and precision for each scenario. This highlights the necessity for refined and flexible strategies in spatio-temporal early prediction.
@@
We address the aforementioned challenges by introducing a spatio-temporal early prediction model based on multi-objective reinforcement learning (STEMO). The key contributions of our study can be summarised as follows:
@@
  • We present a multi-objective reinforcement learning framework designed to optimise both the timeliness and accuracy of spatio-temporal predictions. This approach enhances adaptability to changing patterns through interactive learning from real-time feedback.
@@ 3. Problem Definition
Consider a graph $\mathcal{G}=\{\mathbf{V},\mathbf{E}\}$, where $\mathbf{V}=\{v_{i}\}_{i=1}^{n}$ denotes the set of nodes and $\mathbf{E}=\{e_{ij}\}_{i,j=1}^{n}$ represents the set of edges. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ captures the relationships (e.g., distance) between nodes. Recorded values $\mathbf{X}_{t}=\{x_{t}^{i}\}_{i=1}^{n}$ refer to the measurements obtained from the nodes at time $t$. Predicted values $\widehat{\mathbf{X}}_{T}=\{\hat{x}_{T}^{i}\}_{i=1}^{n}$ are the estimates of the values at time $T$ based on the graph neural network model. For instance, in a traffic speed prediction scenario, $x_{t}^{i}$ could be the speed at sensor $v_{i}$ at time $t$, and $\hat{x}_{T}^{i}$ could be the predicted speed at time $T$. The optimal time series $\mathbf{t^{*}}=\{t^{*}_{i}\}_{i=1}^{n}$ can guide the data observation process. For each node $v_{i}$, the corresponding optimal time $t_{i}^{*}$ is the most appropriate prediction time such that the prediction accuracy is maximised and the time cost minimises. We aim to find the optimal time series that balances prediction accuracy and time cost. The optimal time for node $v_{i}$ is computed using the following expression:
@@
The objective function $\log P(\hat{x}_{T}^{i}|\mathbf{X}_{0:t},\mathcal{G})$ denotes the log-likelihood of the predicted value given the past measurements and the graph structure, which can be seen as a measure of prediction accuracy. The function $\mathrm{Cost}(t)$ represents the time cost of collecting records up to time $t$, and it is assumed to be monotonically increasing as more data is often more expensive to obtain.
@@ 4. Methodology
As shown in Figure 2, the Spatio-Temporal Early Prediction model based on Multi-Objective reinforcement learning (STEMO) consists of three main components: a spatio-temporal predictor, a state generator, and optimal policies for determining the optimal time. In addition, we also introduce how to find hidden preferences.
Figure 2. At time $t$, the encoder processes the recorded values $\mathbf{X}_{0:t}$ to extract spatio-temporal features and generate the hidden state $\mathbf{H}_{t}$. Using $\mathbf{H}_{t}$, the decoder generates a series of predicted values, focusing on $\mathbf{\widehat{X}}^{(t)}_{T}$. The state generator concatenates the node embedding result and $\mathbf{H}_{t}$ to generate the state $\mathbf{s}_{t}$. The policy utilises $\mathbf{s}_{t}$ to determine the optimal time for each node $v_{i}\in\mathbf{V}$ via the action set $\mathbf{a}_{t}=\{a_{t}^{i}\}_{i=1}^{n}$ (halt or wait). ’Wait’ implies that further observation of recorded values is necessary, while ’Halt’ implies that time $t$ is the optimal time $t_{i}^{*}$ for node $v_{i}$, and the corresponding predicted value is recorded in the predicted value $\mathbf{\widehat{X}}_{T}$.
@@ 4.1. The Spatio-temporal Predictor
The main difference between traditional GCN and the Multi-Graph Convolutional Neural network (MGCN) is the addition of multiple time steps in the similarity matrix calculation in the latter. A standard GCN calculates similarity based on spatial proximity or a single time-point temporal similarity, leading to a limitation in capturing time-evolving trends. MGCN improves this by including multiple time-step similarities, capturing more intricate time-evolving correlations, and facilitating early prediction.
@@
Figure 3. Figure (a) shows the time series of nodes $v_{i}$ and $v_{j}$. The dotted line represents the series after the optimal time, and we only need to observe the solid line to make prediction. The two solid dots in Figure (a) correspond to the two red circles in Figure (b). Take the red circle below as an example, it corresponds to the DTW distance between $x_{0:t}^{i}$ and $x_{0:t}^{j}$, which is calculated along the green grid path. We anticipate that node $i$ will acquire feature $x_{t_{j}^{*}}^{j}$ at time $t$ through MGCN, allowing it to make more precise predictions earlier.
Specifically, GCN collects equal-length time series of each node, applies the DTW algorithm to each pair of time series data, and uses dynamic time warping (DTW) distance (Müller, 2007) to create a similarity matrix. DTW is a method that finds an optimal match between two given sequences (e.g., time series) with certain restrictions, which is particularly useful in our setting where we seek to find correlations in temporal data. However, GCN has certain limitations. One limitation is that it does not consider the influence of other nodes before time $t$. As shown in Figure 3, $t$ represents a specific time, and $t^{\prime}$ is an arbitrary time. If $t^{\prime}<t$ and $\mathrm{DTW}(x_{0:t^{\prime}}^{j},x_{t}^{i})=0$, this may result from node $v_{j}$ changing faster than the node $v_{i}$. In such cases, it would be advantageous for node $v_{i}$ to incorporate the feature of node $v_{j}$ at time $t^{\prime}$ in its prediction, as this would enable it to capture the changing trend better and facilitate early prediction. Moreover, if $t^{\prime}=t_{j}^{*}<t$, this would also allow node $i$ to reach the optimal time more quickly. Another limitation of this approach is that it may produce inaccurate results when $\mathrm{DTW}(x_{t_{j}^{*}}^{j},x_{t_{j}^{*}}^{i})>0$, indicating a weak correlation between nodes $v_{j}$ and $v_{i}$. This may not always be true and could lead to sub-optimal predictions. The MGCN addresses these limitations, which is designed to optimize both prediction accuracy and prediction timing. As depicted in Figure 4, the multiple timesteps similarity matrix allows the model to capture correlations better, thus achieving more accurate prediction. In addition, MGCN can incorporate the fast-changing trends from other nodes to make predictions earlier.
@@
The encoder-decoder structure is utilized for its capacity to capture temporal dependencies and handle variable length sequences, as shown in Figure 2. In contrast to previous work, the variation in the number of prediction units (consisting of an MGCN and a GRU) allows the model to dynamically adjust its focus to different parts of the time sequence. At time $t$, the encoder has $t+1$ prediction units. The recorded values $\mathbf{X}_{t}$ are input into MGCN of the last prediction unit in the encoder and the result of graph convolution process $\mathbf{X^{{}^{\prime}}}_{t}$ is then used as input for the GRU. Combining the hidden state $\mathbf{H}_{t-1}$ of GRU output in the previous unit, the GRU outputs hidden state $\mathbf{H}_{t}$. In practice, in the encoder, only the last prediction unit at each time needs to be calculated. At time $t$, the decoder has $T-t-1$ prediction units, and the last prediction unit outputs the candidate predicted values $\mathbf{\widehat{X}}_{T}^{(t)}$.
$\mathbf{\widehat{X}}^{(t)}_{T}=\mathbf{W}\mathbf{H}_{T}^{(t)}+\mathbf{b}$ represents candidate predicted values at time $T$ predicted at time $t$, where $\mathbf{W}$ and $\mathbf{b}$ are learnable parameters. The spatio-temporal predictor is trained by minimizing the mean absolute error between $\mathbf{\widehat{X}}_{T}$ and $\mathbf{X}_{T}$, where predicted values $\mathbf{\widehat{X}}_{T}=\sum_{t=0}^{T-1}\mathbf{a}_{t}\odot\mathbf{\widehat{X}}^{(t)}_{T}$, and $\odot$ denotes the Hadamard product calculation.
@@ 4.2. State Generator
Biased random walk sampling is used to select a node’s neighbours that contribute to its state. The bias term $\alpha_{pq}(t)$ and the transition probability $\min_{t^{\prime}}\mathcal{A}^{\mathcal{T}}_{t,t^{\prime}}$ control the sampling process, favouring nodes that are closer (in terms of reaching their optimal time) and thus more relevant to the early prediction task. Given the source node $v_{i}$ and the current node $v_{j}$, the probability of accessing the next node $v_{k}$ is represented by the bias term $\alpha_{pq}(t)$ multiplied by the transition probability $\mathop{\min}\limits_{t^{\prime}}\mathcal{A}^{\mathcal{T}}_{t,t^{\prime}}$, where $\alpha$ represent bias, and is calculated as follows:

### v3

# STEMO: Early Spatio-temporal Forecasting with Multi-Objective Reinforcement Learning

@@ Abstract.
Accuracy and timeliness are indeed often conflicting goals in prediction tasks. Premature predictions may yield a higher rate of false alarms, whereas delaying predictions to gather more information can render them too late to be useful. In applications such as wildfires, crimes, and traffic jams, timely forecasting are vital for safeguarding human life and property. Consequently, finding a balance between accuracy and timeliness is crucial. In this paper, we propose an early spatio-temporal forecasting model based on Multi-Objective reinforcement learning that can either implement an optimal policy given a preference or infer the preference based on a small number of samples. The model addresses two primary challenges: 1) enhancing the accuracy of early forecasting and 2) providing the optimal policy for determining the most suitable prediction time for each area. Our method demonstrates superior performance on three large-scale real-world datasets, surpassing existing methods in early spatio-temporal forecasting tasks.
Spatio-Temporal Data, Graph Neural Network, Early Detection, Responsible AI
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671922ccs: Computing methodologies Neural networksccs: Computing methodologies Spatial and physical reasoningccs: Applied computing Forecasting
@@ 1. Introduction
Figure 1. Example of three methods for early spatio-temporal forecasting. Each circle is a node (e.g. sensor) with recorded value (such as speed) over time. The blue plane represents the prediction time, and the recorded values after the blue plane are not used for forecasting. The adaptive early forecasting method adjusts data usage and dynamically determines the prediction time for different nodes. The node colour at time $T$ indicates the accuracy, green indicates that the predicted value matches the ground truth, and red indicates otherwise.
@@
Early spatio-temporal forecasting focuses on anticipating future events by examining patterns that change over space and time as early as possible, which is essential to many real-world applications such as epidemiology, environmental studies, and public safety, where early detection can lead to better management or prevention outcomes. For instance, within the context of forecasting the spread of diseases, the timeliness forecasting is considerably more critical than their accuracy. Predictions that are accurate but delayed can lead to the unnecessary loss of thousands of lives. On the other hand, a prediction that is approximately accurate but made in a timely manner can significantly reduce the impact, saving lives.
@@
Figure 1 shows existing approaches to forecast spatio-temporal events are categorised into three distinct strategies: Traditional, Fixed, and Adaptive. For example, Hochreiter et al. (Hochreiter and Schmidhuber, 1997) propose a traditional method that considers the entire recorded data before the forecasting. However, this approach can be time-consuming, particularly with large datasets, and its requirement for complete prior data collection often results in delays, which is unsuitable for time-sensitive applications. Conversely, Li’s approach operates within a set observation period (Li et al., 2017), facilitating fixed timely forecasts within a specific window. Although this strategy enhances promptness, its fixed timing can result in inaccuracies if the observation window is not optimally selected. Moreover, its rigidity in adjusting to temporal data variations and the risk of selection bias in determining the observation window could compromise its precision. The third approach, named adaptive early forecasting, dynamically adjusts the forecasting time based on data characteristics. This idea is inspired by Hartvigsen, who applied this idea in the time-series prediction area (Hartvigsen et al., 2019). This innovative method aims to strike an optimal balance between forecasting accuracy and timeliness, thereby attempting to optimise both the timeliness and accuracy. Its dynamic adaptability provides a significant edge in handling changes over time.
@@
Although Hartivigsen has tried to apply the concept of early prediction in time-series area (Hartvigsen et al., 2019) and achieved some good results in the medical diagnosis, many challenges remain, especially in the spatio-temporal forecasting. (1) Real-time Bi-objective Balance in Highly Complex Environment: Spatio-temporal data demand dynamic solutions, being more complex and subject to rapid changes than time-series data. Traditional multi-objective optimisation methods, which are static and computationally intensive, fail to adapt efficiently across different environments. At the same time, early forecasting necessitates immediate results, as even a one-minute delay can lead to unfavourable outcomes. (2) Spatio-temporal Comprehensive Dependency: Relying solely on distance correlation for spatio-temporal forecasting may fall short in quickly capturing essential data characteristics. Although distance correlation effectively reflects spatial connections among nodes, it overlooks the crucial temporal aspects embedded within the data, which are vital for precise forecasting. (3) Hidden Preferences Discovery: The equilibrium between timeliness and accuracy varies across tasks, revealing diverse preferences. Identifying these subtle preferences is a complex challenge that demands a profound comprehension of the goals at hand and formulating an ideal balance between promptness and precision for each scenario. This highlights the necessity for refined and flexible strategies in early spatio-temporal forecasting.
@@
We address the aforementioned challenges by introducing a early spatio-temporal forecasting model based on multi-objective reinforcement learning (STEMO). The key contributions of our study can be summarised as follows:
@@
  • We present a multi-objective reinforcement learning framework designed to optimise both the timeliness and accuracy of spatio-temporal forecasting. This approach enhances adaptability to changing patterns through interactive learning from real-time feedback.
@@ 3. Problem Definition
Consider a graph $\mathcal{G}=\{\mathbf{V},\mathbf{E}\}$, where $\mathbf{V}=\{v_{i}\}_{i=1}^{n}$ denotes the set of nodes and $\mathbf{E}=\{e_{ij}\}_{i,j=1}^{n}$ represents the set of edges. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ captures the relationships (e.g., distance) between nodes. Recorded values $\mathbf{X}_{t}=\{x_{t}^{i}\}_{i=1}^{n}$ refer to the measurements obtained from the nodes at time $t$. forecasted values $\widehat{\mathbf{X}}_{T}=\{\hat{x}_{T}^{i}\}_{i=1}^{n}$ are the estimates of the values at time $T$ based on the graph neural network model. For instance, in a traffic speed forecasting scenario, $x_{t}^{i}$ could be the speed at sensor $v_{i}$ at time $t$, and $\hat{x}_{T}^{i}$ could be the forecasted speed at time $T$. The optimal time series $\mathbf{t^{*}}=\{t^{*}_{i}\}_{i=1}^{n}$ can guide the data observation process. For each node $v_{i}$, the corresponding optimal time $t_{i}^{*}$ is the most appropriate that the forecasting accuracy is maximised and the time cost minimises. We aim to find the optimal time series that balances forecasting accuracy and time cost. The optimal time for node $v_{i}$ is computed using the following expression:
@@
The objective function $\log P(\hat{x}_{T}^{i}|\mathbf{X}_{0:t},\mathcal{G})$ denotes the log-likelihood of the forecasted value given the past measurements and the graph structure, which can be seen as a measure of the accuracy. The function $\mathrm{Cost}(t)$ represents the time cost of collecting records up to time $t$, and it is assumed to be monotonically increasing as more data is often more expensive to obtain.
@@ 4. Methodology
As shown in Figure 2, the Early Spatio-Temporal Forecasting model based on Multi-Objective reinforcement learning (STEMO) consists of three main components: a spatio-temporal predictor, a state generator, and optimal policies for determining the optimal time. In addition, we also introduce how to find hidden preferences.
Figure 2. At time $t$, the encoder processes the recorded values $\mathbf{X}_{0:t}$ to extract spatio-temporal features and generate the hidden state $\mathbf{H}_{t}$. Using $\mathbf{H}_{t}$, the decoder generates a series of forecasted values, focusing on $\mathbf{\widehat{X}}^{(t)}_{T}$. The state generator concatenates the node embedding result and $\mathbf{H}_{t}$ to generate the state $\mathbf{s}_{t}$. The policy utilises $\mathbf{s}_{t}$ to determine the optimal time for each node $v_{i}\in\mathbf{V}$ via the action set $\mathbf{a}_{t}=\{a_{t}^{i}\}_{i=1}^{n}$ (halt or wait). ’Wait’ implies that further observation of recorded values is necessary, while ’Halt’ implies that time $t$ is the optimal time $t_{i}^{*}$ for node $v_{i}$, and the corresponding forecasted value is recorded in $\mathbf{\widehat{X}}_{T}$.
@@ 4.1. The Spatio-temporal Predictor
The main difference between traditional GCN and the Multi-Graph Convolutional Neural network (MGCN) is the addition of multiple time steps in the similarity matrix calculation in the latter. A standard GCN calculates similarity based on spatial proximity or a single time-point temporal similarity, leading to a limitation in capturing time-evolving trends. MGCN improves this by including multiple time-step similarities, capturing more intricate time-evolving correlations, and facilitating timeliness.
@@
Figure 3. Figure (a) shows the time series of nodes $v_{i}$ and $v_{j}$. The dotted line represents the series after the optimal time, and we only need to observe the solid line to make forecasting. The two solid dots in Figure (a) correspond to the two red circles in Figure (b). Take the red circle below as an example, it corresponds to the DTW distance between $x_{0:t}^{i}$ and $x_{0:t}^{j}$, which is calculated along the green grid path. We anticipate that node $i$ will acquire feature $x_{t_{j}^{*}}^{j}$ at time $t$ through MGCN, allowing it to make more precise forecasting earlier.
Specifically, GCN collects equal-length time series of each node, applies the DTW algorithm to each pair of time series data, and uses dynamic time warping (DTW) distance (Müller, 2007) to create a similarity matrix. DTW is a method that finds an optimal match between two given sequences (e.g., time series) with certain restrictions, which is particularly useful in our setting where we seek to find correlations in temporal data. However, GCN has certain limitations. One limitation is that it does not consider the influence of other nodes before time $t$. As shown in Figure 3, $t$ represents a specific time, and $t^{\prime}$ is an arbitrary time. If $t^{\prime}<t$ and $\mathrm{DTW}(x_{0:t^{\prime}}^{j},x_{t}^{i})=0$, this may result from node $v_{j}$ changing faster than the node $v_{i}$. In such cases, it would be advantageous for node $v_{i}$ to incorporate the feature of node $v_{j}$ at time $t^{\prime}$, this would enable it to capture the changing trend better and facilitate early forecasting. Moreover, if $t^{\prime}=t_{j}^{*}<t$, this would also allow node $i$ to reach the optimal time more quickly. Another limitation of this approach is that it may produce inaccurate results when $\mathrm{DTW}(x_{t_{j}^{*}}^{j},x_{t_{j}^{*}}^{i})>0$, indicating a weak correlation between nodes $v_{j}$ and $v_{i}$. This may not always be true and could lead to sub-optimal outcome. The MGCN addresses these limitations, which is designed to optimize both forecasting accuracy and timeliness. As depicted in Figure 4, the multiple timesteps similarity matrix allows the model to capture correlations better, thus achieving more accurate outcome. In addition, MGCN can incorporate the fast-changing trends from other nodes to make forecasting earlier.
@@
The encoder-decoder structure is utilized for its capacity to capture temporal dependencies and handle variable length sequences, as shown in Figure 2. In contrast to previous work, the variation in the number of units (consisting of an MGCN and a GRU) allows the model to dynamically adjust its focus to different parts of the time sequence. At time $t$, the encoder has $t+1$ units. The recorded values $\mathbf{X}_{t}$ are input into MGCN in the unit $t+1$, and the result of graph convolution process $\mathbf{X^{{}^{\prime}}}_{t}$ is then used as input for the GRU. Combining the hidden state $\mathbf{H}_{t-1}$ of GRU output in the previous unit, the GRU outputs hidden state $\mathbf{H}_{t}$. In practice, in the encoder, only the last prediction unit at each time needs to be calculated. At time $t$, the decoder has $T-t-1$ units, and the last unit outputs the candidate forecasting values $\mathbf{\widehat{X}}_{T}^{(t)}$.
$\mathbf{\widehat{X}}^{(t)}_{T}=\mathbf{W}\mathbf{H}_{T}^{(t)}+\mathbf{b}$ represents candidate forecasted values at time $T$ forecasted at time $t$, where $\mathbf{W}$ and $\mathbf{b}$ are learnable parameters. The spatio-temporal predictor is trained by minimizing the mean absolute error between $\mathbf{\widehat{X}}_{T}$ and $\mathbf{X}_{T}$, where forecasted values $\mathbf{\widehat{X}}_{T}=\sum_{t=0}^{T-1}\mathbf{a}_{t}\odot\mathbf{\widehat{X}}^{(t)}_{T}$, and $\odot$ denotes the Hadamard product calculation.
@@ 4.2. State Generator
Biased random walk sampling is used to select a node’s neighbours that contribute to its state. The bias term $\alpha_{pq}(t)$ and the transition probability $\min_{t^{\prime}}\mathcal{A}^{\mathcal{T}}_{t,t^{\prime}}$ control the sampling process, favouring nodes that are closer (in terms of reaching their optimal time) and thus more relevant to the early forecasting task. Given the source node $v_{i}$ and the current node $v_{j}$, the probability of accessing the next node $v_{k}$ is represented by the bias term $\alpha_{pq}(t)$ multiplied by the transition probability $\mathop{\min}\limits_{t^{\prime}}\mathcal{A}^{\mathcal{T}}_{t,t^{\prime}}$, where $\alpha$ represent bias, and is calculated as follows: